OpenStudy (anonymous):

solve the equation. x+1/3=-6/3, for the given variable. write answer as a reduced fraction.

OpenStudy (anonymous):

I don't really understand the question, I think it's asking to solve for X? In that case, x=-7/3

OpenStudy (anonymous):

it was correct!! thank you :)

OpenStudy (anonymous):

my guess is \[\huge \frac{x+1}{3}=\frac{-6}{3}\] but i could be wrong

OpenStudy (anonymous):

No problem! If you need help with more problems just tag me.